OpenStudy (anonymous):

Aran moved a cone-shaped pile of sand that had a height of 6 ft and a radius of 3 ft. He used all of the sand to fill a cylindrical pit with a radius of 6 ft. How high did the sand reach in the pit? Use 3.14 to approximate pi and express your final answer in tenths. ft

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the volume of the cone?

OpenStudy (luffingsails):

So, first calculate the volume of sand in the cone. Then set that volume equal to the equation of the volume of a cylinder with radius 6 ft... then solve for the height of the cylinder.

OpenStudy (anonymous):

it doesn't say @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use the formula given on the previous problem to calculate the volume of the cone

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i confused my self can you set up the equation for me?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large V = \frac{1}{3}*\pi*r^2*h\] is the formula you need

jimthompson5910 (jim_thompson5910):

for the volume of a cone

OpenStudy (anonymous):

what should i do with 2 radius though

jimthompson5910 (jim_thompson5910):

"Aran moved a cone-shaped pile of sand that had a height of 6 ft and a radius of 3 ft." r = 3 h = 6

OpenStudy (anonymous):

ok so it would be 3^2*6*3.14?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is that equal to

OpenStudy (anonymous):

ok so it is equal to 169.56

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Volume of a Cylinder \[\Large V = \pi*r^2*h\]

OpenStudy (anonymous):

so that is correct?

jimthompson5910 (jim_thompson5910):

plug V = 169.56 into \[\Large V = \pi*r^2*h\] along with r = 6

jimthompson5910 (jim_thompson5910):

then solve for h

OpenStudy (anonymous):

339.12?

OpenStudy (anonymous):

or is it 113.04

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large V = \pi*r^2*h\] \[\Large 169.56 = 3.14*6^2*h\] \[\Large h = ???\]

OpenStudy (anonymous):

idk what to do

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have to solve for h

jimthompson5910 (jim_thompson5910):

look at the last steps I posted above

OpenStudy (anonymous):

h=0.870

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is 3.14*6^2 equal to?

OpenStudy (anonymous):

113.04 i kept getting that earlier but i thought it was wrong

jimthompson5910 (jim_thompson5910):

now you have 169.56 = 113.04*h

jimthompson5910 (jim_thompson5910):

divide both sides in 169.56 = 113.04*h by 113.04 to isolate h

OpenStudy (anonymous):

SO 113.04*113.04

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

169.56 = 113.04*h 169.56/113.04 = h h = ???

OpenStudy (anonymous):

oh ok 1.5

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so thats my answer?

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

@jim_thompson5910 was that my answer?